OpenStudy (anonymous):

SHORT CUT FOR SOLVING 8*88*888*8888

OpenStudy (gorv):

these are in multiplication??? are u sure

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i know the shortcut for 8+88+888... but i don't know how to solve this series in multiplication within seconds.reply me with a solution please!

OpenStudy (anonymous):

do you know the shortcut for this question?

OpenStudy (gorv):

is it a Q really??

OpenStudy (gorv):

i think u made it yourself u have to mulyiply it .....if i will get a solution i will inform you it will rach upto infinity loll

OpenStudy (liamschumm):

Factor the 8 out. \[8∗88∗888∗8888=8∗(1∗11∗111∗1111)=8∗(11∗111∗1111).\]

OpenStudy (liamschumm):

Oh my, the equation editor is not happy. This is what I meant to post: 8∗88∗888∗8888=8∗(1∗11∗111∗1111)=8∗(11∗111∗1111)

OpenStudy (liamschumm):

Oh my, again. One more try. 8*88*888*8888=8(1*11*111*1111)=8(11*111*1111).

OpenStudy (liamschumm):

There we go.